OpenStudy (anonymous):

4(2x+11)=76 ASAP PLEASE!

OpenStudy (happy_to_help):

First thing you do is distribute. so 4times2x=8x 4times11=44 8x+44=76 Subtract 44 from both sides 8x=32 32 divided by 8 =4 x=4

OpenStudy (happy_to_help):

Do you understand?

OpenStudy (anonymous):

Can you help me with another problem Please

OpenStudy (happy_to_help):

Sure

OpenStudy (happy_to_help):

Whats the question?

OpenStudy (anonymous):

14(x+1)=-7(4+x)

OpenStudy (happy_to_help):

First start by doing one side. Distribute 14timesX=14x 14times1=14 14x+14 -7times4=-28 -7timesX=-7x -28+(-7x)= 14x+14=-7-28 Add 7 to both sides 21x+14=-28 Subtract 14 from both sides 21x=-42 -42dividedby21=-2 x=-2

OpenStudy (happy_to_help):

Do you understand?

OpenStudy (happy_to_help):

Need anymore help?

OpenStudy (anonymous):

Nope but maybe tommorow but sorry about that Thank you sm

OpenStudy (happy_to_help):

No promblem. Any time